OpenStudy (anonymous):

How would you prove that the limit dos not exist? I'll include the equation in the answers

OpenStudy (anonymous):

\[\lim (x,y)->(0,0) \frac{ y }{ \sqrt{x ^{2}-y ^{2}} }\]

OpenStudy (anonymous):

Using the two-path test

OpenStudy (turingtest):

try going along the line x=0

OpenStudy (turingtest):

or rather y=0

OpenStudy (anonymous):

ok, what would the opposite direction be? If you go x = 0 you get an imaginary number

OpenStudy (turingtest):

yeah that was my mistake, but going along y=0 the limit is...?

OpenStudy (anonymous):

The limit is zero

OpenStudy (turingtest):

right, now try something that is likely not to give out all zeros, for example along the line y=x-1

OpenStudy (anonymous):

ok that worked, thank you!

OpenStudy (turingtest):

welcome :)